(i) Use method of separation of variables in rectangular coordinates to obtain the following harmonic functions which are bounded in the upper half plane y > 0 of R2,
e−λy cosλx, e−λy sinλx; λ ≥0.
(ii) By integration with respect to the parameter λ, obtain the functions
x / x2 +y2, y / x2 +y2 which are harmonic in the upper half plane y > 0 of R2.
1
Expert's answer
2020-09-07T17:30:39-0400

i) Let the required harmonic function be F such that,

F(x,y)=X(x,t).Y(y,t).....(1)

As per the question the boundary condition are y>0,(0,0),to "e^{-\lambda y}Cos\lambda x,e^{-\lambda y}{Sin\lambda x}"


from equation 1..

we have

"\frac{ d^2X}{ dx^2}+\frac{d^2Y}{dy^2}=0"

The above equation can be written according to separation of variable

"\frac{ d^2X}{ Xdx^2}=-\frac{d^2Y}{Y dy^2}" = k"^2"

"\frac{d^2X}{dx^2}=k^2x, \frac{d^2Y}{dy^2}=-k^2y"

The solution of the above equation can be written as

X(x,t)="c_1coskx+c_2sinkx,Y(y,t)=c_3e^{kx}+c_4e^{ky}"


Then the harmonic function is given by,

"F(x,y)=(c_1coskx+c_2sinkx)(c_3e^{kx}+c_4e^{ky})"

This is the required function.

ii)As "x=e^{-\lambda y}Cos\lambda x"

y="x=e^{-\lambda y}Sin\lambda x"

then"\frac{x^2}{x^2+y^2}=\frac{e^{-2\lambda y} Cos^2\lambda x}{e^{-2\lambda y} Cos^2\lambda x+e^{-2\lambda y} Sin^2\lambda x}"

"\frac{x^2}{x^2+y^2}=\frac{e^{-2\lambda y} Cos^2\lambda x}{e^{-2\lambda y} (Cos^2\lambda x+ Sin^2\lambda x)}"


"\frac{x^2}{x^2+y^2}= Cos^2\lambda x"


similiarly"\frac{y^2}{x^2+y^2}=Sin^2\lambda x"

Now the integrate both of the function with respect to "\lambda"


"\int\frac{x^2}{x^2+y^2}d\lambda=\int Cos^2\lambda d\lambda=\int (\frac{1+Cos2\lambda x}{2})d\lambda"

="\frac{\lambda +\frac{Sin2\lambda x}{2x}}{2}"

Now integrate

"\int\frac{y^2}{x^2+y^2}d\lambda=\int Sin^2\lambda d\lambda"

="\int (\frac{1-Cos2\lambda x}{2})d\lambda"


="\frac{\lambda -\frac{Sin2\lambda x}{2x}}{2}"

These are the required functions.